OpenStudy (anonymous):

int_{}^{}\frac{ 9 }{ 2+sinx }dx

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 9 }{ 2+sinx }dx\]

OpenStudy (anonymous):

substitute t= tan(x/2) then sin x = 2t/(1+t^2) dx = 1/(1+t^2)

OpenStudy (anonymous):

reference : http://openstudy.com/users/hartnn#/updates/50960518e4b0d0275a3ccfba

OpenStudy (anonymous):

see first tip there....

OpenStudy (anonymous):

can u solve that now ?

OpenStudy (anonymous):

i'm stuck with \[9\int\limits_{}^{}\frac{ dt }{ t ^{2}+t+1 }\]

OpenStudy (anonymous):

know the method of completing the square ? u need to complete the square in deniminator

OpenStudy (anonymous):

and then use 1/(x^2+a^2) formula formula 16 from that list

OpenStudy (anonymous):

yup it's \[(t ^{2}+\frac{ 1 }{ 2})^{2}+\frac{ 3 }{ 4 }\] right?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

now its in the form x^2+a^2 or u can also put u=t+1/2

OpenStudy (anonymous):

*t

OpenStudy (anonymous):

its actually this : \((t +\frac{ 1 }{ 2})^{2}+\frac{ 3 }{ 4 }\)

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

thanks got it. i'm trying the partial fractions but didn't get it, is it possible w/ partial fractions?

OpenStudy (anonymous):

u can't factorize denominator, so no....

OpenStudy (anonymous):

not into real factors.

OpenStudy (anonymous):

right ?

OpenStudy (anonymous):

ok thanks